# Layered Division and Global Allocation for Community Detection in Multilayer Network

### v1

@@ I Introduction
Community detection, the task of identifying groups of nodes that are densely connected and share structural similarities in a network, has long been a central focus of web mining and network science[fortunato2010community]. It supports many applications in diverse domains, like biology, economics, and sociology[javed2018community]. Traditional methods, such as modularity maximization[GM], label propagation[LabelPropagation], or graph embedding[grover2016node2vec], have achieved lots of successes for community detection in single-layer networks, where only one type of relation is modeled and all edges are hence of the same type.
@@
In reality, entities in Web systems and other domains often have multiple yet different types of relations. For example, users may be connected simultaneously through offline friendships, online hyperlinks, or financial transactions. Collapsing these heterogeneous interactions into a single layer discards crucial relational information and risks obscuring the true community structure. Multilayer networks address this limitation by preserving all relational types, each by one network layer with edges representing one type of relation. Detecting communities in multilayer networks has numerous applications, such as uncovering the social groups of users on different online platforms [boccaletti2014structure].
Recently, many neural network-based solutions have been proposed for the community detection in multilayer networks [NFACC, MDeepWalk, MCLCD, NACC, GEAM, DMGI, HDMI]. However, they all follow a kind of representation fusion and global division (RFGD) paradigm: neural encoders are designed to first learn layer-wise nodes’ representations, that is, a representation is learned in each network layer for each node. Next, fusion mechanisms are designed to fuse layer-wise representations into one global representation, that is, a global representation is fused for each node. Finally, clustering algorithms are designed to divide nodes into communities based on computing similarities between their global representations.
The RFGD paradigm has evolved significantly. Early strategies, as shown in Fig. 1(a), relied on network flattening , which collapses the multilayer structure into a single graph [berlingerio2013abacus], so that standard detection algorithms (e.g., Label Propagation [LabelPropagation], Louvain [Louvain] and Greedy Modularity [GM]) can be applied. In contrast, modern neural solutions are illustrated in Figure 1(b). For example, MGCCN [MGCCN] evaluates the relevance coefficients of neighboring nodes for learning layer-wise representations; In order to fuse inter-layer semantic information, GEAM [GEAM] first applied a self-attention layer to learn layer-aware representations, then utilized an adaptive fusion mechanism to obtain fused representations. Then, the $K$-means algorithm is applied to cluster nodes into communities based on the fused representations.
@@
In this paper, we argue that such an RFGD paradigm might not be the best choice for the community detection in multilayer networks (CDMN) task. It suffers from three key limitations. (1) Layer fidelity. Representation fusion may blur the distinctions of individual layers, either collapsing them into a single view or mixing them with concatenation/additive attention, which weakens the ability to attribute communities to specific relation types. (2) Representation expressiveness. Most recent approaches rely on the graph neural networks (e.g., GCN [kipf2016semi]/GAT [velickovic2017graph]) for per-layer encoding, but these models are prone to limited expressivity [xu2018powerful, egressy2024provably], over-smoothing [li2018deeper, song2023ordered], and over-squashing [alon2020bottleneck]; as a result, they struggle to capture higher-order structural and long-range topological characteristics on each network layer (e.g., cycles [chen2020can, chen2019equivalence]). More expressive, attention-based encoders (e.g., transformers [maron2018invariant, kim2021transformers]) offer a stronger basis for layer-wise learning. (3) Objective mismatch. Most methods optimize a kind of objective functions for reconstruction or contrastive surrogates, rather than the actual goal of community detection, leaving representation learning and community division decoupled. Together, these challenges call for a new paradigm that is more layer-faithful and representation-expressive, as well as a more task-aligned training objective for the CDMN task.
@@ II Related Works
Information Fusion and Community Detection in multi-layer Networks. A central challenge in multi-layer network analysis is integrating information across layers without losing critical structural details. Methodologies have evolved from direct structural aggregation to more nuanced, learning-based approaches. The most direct strategies involve ”flattening” or aggregating the multi-layer graph into a single, weighted graph, upon which standard community detection algorithms can be applied [fortunato2010community]. For example, ABACUS [berlingerio2013abacus] sums the adjacency matrices of each layer; Cozzo et al. [cozzo2016multilayer] proposed a method to construct a larger ”supra-graph” that represents node-layer pairs. While computationally simple, these methods share a fundamental limitation: the premature compression of information via network flattening inevitably discards the unique topological details of each layer, potentially misleading the global community divisions.
@@
A more sophisticated RFGD paradigm has emerged for the CDMN task with the rise of graph neural networks (GNNs) [kipf2016semi]. The general pipeline involves first learning layer-wise embeddings using methods like DeepWalk [perozzi2014deepwalk] and subsequently fusing them to produce a unified representation. In [GCFM], a graph convolutional auto-encoder is proposed for learning a low-dimensional vector encoding for each node before clustering. More advanced models use attention mechanisms or network augmentation to learn a weighted combination of layer embeddings, as seen in models like GEAM[GEAM], NFACC[NFACC], and Kazim et al. [kazim2025multilayer]. However, even these sophisticated methods ultimately merge distinct layer representations into a single global representation for each node. This final fusion step can blur heterogeneous relational signals, losing the strict separation of layer-wise topologies.
@@
Learning Paradigms and Objective Functions for Graph Clustering. Beyond the fusion strategy, the learning objective function profoundly impacts the quality of the detected communities. The majority of recent deep learning methods follow a two-stage pipeline that separates representation learning from the final clustering step. In the first stage, a GNN encoder learns node embeddings in a self-supervised manner, often using contrastive learning frameworks like Deep Graph Infomax (DGI) [velivckovic2018deep], DMGI [DMGI], and NACC[NACC]. In the second stage, a standard clustering algorithm like K-means is applied to cluster these embeddings. This decoupled pipeline, however, creates a significant objective mismatch. The GNN is trained to optimize a surrogate objective, such as a contrastive loss, which is not a direct measure of the clustering quality. An embedding optimized for general representation quality is not guaranteed to be optimal for partitioning the graph into dense communities.
@@
To resolve this mismatch, an end-to-end paradigm has emerged that directly optimizes a clustering-specific objective function. For single-layer networks, the recent work [tsitsulin2023graph] introduced a fully differentiable version of the classic modularity metric, allowing a GNN to be trained to directly maximize the clustering quality. Our proposed LDGA framework extends this powerful insight to multi-layer networks. By developing a differentiable multi-layer modularity loss, we align the entire learning process with the goal of community detection in multilayer networks, unifying feature learning and graph clustering into an end-to-end trainable system.
@@ IV-A Node Feature Initialization
Our framework supports multilayer graphs with optional node attributes. If raw node features are provided, they can be directly incorporated as input to each layer. If no attributes are available, which is common in many benchmark datasets, we initialize features via an unsupervised embedding method such as DeepWalk [perozzi2014deepwalk], applied independently to each layer. This produces per-layer embeddings ${\bm{Z}}^{(s)}=\left[{\bm{z}}^{(s)}_{1},{\bm{z}}^{(s)}_{2},\dots,{\bm{z}}^{(s)}_{n}\right]^{T}\in\mathbb{R}^{N\times d}$ that capture local connectivity patterns while preserving layer identity. Stacking them yields a feature tensor ${\bm{Z}}\in\mathbb{R}^{L\times N\times d}$.
@@ IV-B Community-Latent Encoder (CLE)
Most existing deep learning methods focus on designing neural network architectures from the perspective of nodes and network layers [teng2024multi, GEAM, liu2024motif], while overlooking community-level semantics. To enrich representations, we introduce a community-latent encoder (CLE), which enriches node features with learnable prototypes that represent latent communities. For each layer $s$, CLE maintains a learnable prototype matrix ${\bm{E}}^{(s)}\in\mathbb{R}^{\kappa\times d}$, where each row vector represents a learnable latent prototype for one community.
@@ IV-E Multi-layer Modularity Maximization Loss
Inspired by Tsitsulin et al.[tsitsulin2023graph], who introduced a differentiable relaxation of the Newman–Girvan modularity, we extend this idea to multilayer network. We propose a loss function, ${\mathcal{L}}_{\text{soft}}$, which couples differentiable multilayer modularity with a normalized cluster balance regularizer:
@@ V-A Experiment Setup
To evaluate the performance of our proposed method, we conduct experiments on both synthetic and real-world datasets. Synthetic datasets mainly refer to mLFR [mLFR], which is a baseline model for generating multilayer networks. The mixing parameter $\mu$ in the mLFR datasets, indicating the probability of edges connecting to the nodes of the other communities, can be adjusted. A smaller $\mu$ indicates more obvious community structures, while a larger $\mu$ indicates more vague community structures. Table I presents the detailed parameter settings of mLFR datasets in our experiment.
We also select ten extensive real-world multilayer networks datasets from domains covering society (AUCs [AUCs], CKM [CKM], Lazega [lazega2001collegial, lazega2006], RM [RM], Vickers [vickers1981representing]), biology (C.Elegans [CElegans], Drosophila [Drosophila, SACCH], SACCHPOMB [Drosophila, SACCH]), finance Kapferer [kapferer]), transportation (EUAir [EUair]). They are all collected manually and without ground-truth labels. Table II presents their basic information.
@@
In order to verify the performance of our method, we select nine state-of-the-art community detection methods, which can be divided into the following categories. The first group is traditional community detection algorithms: GenLouvain(GL) [GenLouvain], M-Infomap(MI) [MInfo], MDLPA [MDLPA], MolTi [MolTi]. The second group is five multiplex embedding algorithms: HDMI [HDMI], DMGI [DMGI], GEAM [GEAM], NFACC [NFACC], NACC [NACC].
@@
We use common evaluation metrics of community detection to evaluate the performance of all the methods, including NMI (Normalized Mutual Information) [NMI], ARI (Adjusted Rand Index) [ARI], Purity [Purity], and multilayer modularity [mucha2010community]. The maximum value of multilayer modularity is 1, and a higher modularity means a clearer community structure. NMI, ARI, and Purity are only applied to synthetic datasets since they need ground-truth labels, while the modularity of detected communities is applied to both synthetic and real-world datasets. The resolution parameter and coupling parameter of modularity are both set to 1. For all experiments, we run 10 trials with different random seeds, and report the best results in Tables III, IV, and Fig. 3.
@@ VI Conclusion
[Additional Experiment Details]
Dataset Details
In this section, we present the detailed description of the synthetic and real-world dataset used in this paper. The synthetic datasets mainly refer to mLFR[mLFR], which is a baseline model for generating multilayer networks. We also select ten common multilayer networks to conduct experiment. They are all collected manually and without ground-truth labels.
  • mLFR[mLFR]: We mainly adjust the mixing parameter $\mu$ in mLFR datasets, which is the probability of edges connecting to the nodes of the other communities. A smaller $\mu$ indicates more obvious community structures, while a larger $\mu$ indicates more vague community structures. Table I presents the detailed parameter settings of mLFR datasets in our experiment.
The ten real-world datasets were carefully chosen from domains like society, biology, finance, commerce, and others.
  • AUCs[AUCs]: This multiplex social network contains five types of online and offline relationships (e.g., Facebook, Work, Lunch) among employees at a university’s Computer Science department.
  • C.Elegans[CElegans]: This is the connectome of the Caenorhabditis elegans worm, where layers represent different types of synaptic junctions, such as electric and chemical connections.
  • CKM[CKM]: This dataset contains relationships among physicians in four towns in Illinois, originally collected to study how network ties influenced the adoption of a new drug.
  • Drosophila[Drosophila][SACCH]: This network represents different types of genetic interactions for organisms, sourced from the Biological General Repository for Interaction Datasets (BioGRID).
  • EUAir[EUair]: This is a transportation network composed of 37 layers, with each layer corresponding to a different airline operating in Europe.
  • Kapferer[kapferer]: This dataset documents interactions in a tailor shop in Zambia, where layers represent two different types of interactions recorded seven months apart.
  • Lazega[lazega2001collegial][lazega2006]: This multiplex social network consists of 3 kinds of interactions(Co-work, Friendship and Advice) between partners and associates of a corporate law partnership.
  • RM[RM]: This dataset records three social relationships (friendship and proximity at/outside work) between students at the MIT Media Laboratory and MIT Sloan business school.
  • SACCHPOMB[Drosophila][SACCH]: Sourced from the BioGRID repository, this network considers various types of genetic interactions for the fission yeast organism.
  • Vickers[vickers1981representing]: This social network data was collected from 29 seventh-grade students who were asked to nominate classmates based on several different relationships.
Baseline Details
In this section, we present a detailed description of the selected nine state-of-the-art community detection methods, which can be divided into the following categories. The first group is traditional community detection algorithms:
  • GenLouvain(GL)[GenLouvain]: A classic method that detects communities by maximizing modularity across layers, treating multilayer structure as a supra-adjacency matrix.
  • M-Infomap(MI)[MInfo]: A method based on the ”Map Equation,” which identifies communities by minimizing the description length of random walks.
  • MDLPA[MDLPA]: A multi-dimensional label propagation algorithm that extends the traditional method by propagating labels alternately between layers to jointly optimize the community partition.
  • MolTi[MolTi]: This algorithm identifies communities by measuring the consistency between network layers, then performs consensus clustering on the resulting similarity matrices.
The second group is five multiplex embedding algorithms.
  • HDMI[HDMI]: A high-order deep learning method that captures complex structural features in multilayer networks and combines them with deep transfer learning for community detection.
  • DMGI[DMGI]: A contrastive learning-based algorithm that extends Deep Graph Infomax to multiplex networks to generate embeddings for community detection and other tasks
  • GEAM[GEAM]: A graph-enhanced attention model that uses an attention mechanism to integrate node attributes and structural information for generating high-quality node embeddings.
  • NFACC[NFACC]: A GCN-based method combines network feature augmentation with contrastive learning to integrate intra-layer and inter-layer information for robust multi-layer community detection.
  • NACC[NACC]: A self-supervised method fuses intra- and inter-layer features to create a learnable ”anchor” network, which is then used in a contrastive learning framework to train the model.

### v2

@@ I Introduction
Community detection, the task of identifying groups of nodes that are densely connected and share structural similarities in a network, has long been a central focus of web mining and network science[22]. It supports many applications in diverse domains, like biology, economics, and sociology[25]. Traditional methods, such as modularity maximization[14], label propagation[41], or graph embedding[24], have achieved lots of successes for community detection in single-layer networks, where only one type of relation is modeled and all edges are hence of the same type.
@@
In reality, entities in Web systems and other domains often have multiple yet different types of relations. For example, users may be connected simultaneously through offline friendships, online hyperlinks, or financial transactions. Collapsing these heterogeneous interactions into a single layer discards crucial relational information and risks obscuring the true community structure. Multilayer networks address this limitation by preserving all relational types, each by one network layer with edges representing one type of relation. Detecting communities in multilayer networks has numerous applications, such as uncovering the social groups of users on different online platforms [5].
Recently, many neural network-based solutions have been proposed for the community detection in multilayer networks [47, 44, 9, 49, 54, 39, 26]. However, they all follow a kind of representation fusion and global division (RFGD) paradigm: neural encoders are designed to first learn layer-wise nodes’ representations, that is, a representation is learned in each network layer for each node. Next, fusion mechanisms are designed to fuse layer-wise representations into one global representation, that is, a global representation is fused for each node. Finally, clustering algorithms are designed to divide nodes into communities based on computing similarities between their global representations.
The RFGD paradigm has evolved significantly. Early strategies, as shown in Fig. 1(a), relied on network flattening , which collapses the multilayer structure into a single graph [3], so that standard detection algorithms (e.g., Label Propagation [41], Louvain [4] and Greedy Modularity [14]) can be applied. In contrast, modern neural solutions are illustrated in Figure 1(b). For example, MGCCN [33] evaluates the relevance coefficients of neighboring nodes for learning layer-wise representations; In order to fuse inter-layer semantic information, GEAM [54] first applied a self-attention layer to learn layer-aware representations, then utilized an adaptive fusion mechanism to obtain fused representations. Then, the $K$-means algorithm is applied to cluster nodes into communities based on the fused representations.
@@
In this paper, we argue that such an RFGD paradigm might not be the best choice for the community detection in multilayer networks (CDMN) task. It suffers from three key limitations. (1) Layer fidelity. Representation fusion may blur the distinctions of individual layers, either collapsing them into a single view or mixing them with concatenation/additive attention, which weakens the ability to attribute communities to specific relation types. (2) Representation expressiveness. Most recent approaches rely on the graph neural networks (e.g., GCN [30]/GAT [51]) for per-layer encoding, but these models are prone to limited expressivity [55, 21], over-smoothing [32, 45], and over-squashing [1]; as a result, they struggle to capture higher-order structural and long-range topological characteristics on each network layer (e.g., cycles [12, 13]). More expressive, attention-based encoders (e.g., transformers [36, 29]) offer a stronger basis for layer-wise learning. (3) Objective mismatch. Most methods optimize a kind of objective functions for reconstruction or contrastive surrogates, rather than the actual goal of community detection, leaving representation learning and community division decoupled. Together, these challenges call for a new paradigm that is more layer-faithful and representation-expressive, as well as a more task-aligned training objective for the CDMN task.
@@ II Related Works
Information Fusion and Community Detection in multi-layer Networks. A central challenge in multi-layer network analysis is integrating information across layers without losing critical structural details. Methodologies have evolved from direct structural aggregation to more nuanced, learning-based approaches. The most direct strategies involve ”flattening” or aggregating the multi-layer graph into a single, weighted graph, upon which standard community detection algorithms can be applied [22]. For example, ABACUS [3] sums the adjacency matrices of each layer; Cozzo et al. [16] proposed a method to construct a larger ”supra-graph” that represents node-layer pairs. While computationally simple, these methods share a fundamental limitation: the premature compression of information via network flattening inevitably discards the unique topological details of each layer, potentially misleading the global community divisions.
@@
A more sophisticated RFGD paradigm has emerged for the CDMN task with the rise of graph neural networks (GNNs) [30]. The general pipeline involves first learning layer-wise embeddings using methods like DeepWalk [40] and subsequently fusing them to produce a unified representation. In [8], a graph convolutional auto-encoder is proposed for learning a low-dimensional vector encoding for each node before clustering. More advanced models use attention mechanisms or network augmentation to learn a weighted combination of layer embeddings, as seen in models like GEAM[54], NFACC[47], and Kazim et al. [28]. However, even these sophisticated methods ultimately merge distinct layer representations into a single global representation for each node. This final fusion step can blur heterogeneous relational signals, losing the strict separation of layer-wise topologies.
@@
Learning Paradigms and Objective Functions for Graph Clustering. Beyond the fusion strategy, the learning objective function profoundly impacts the quality of the detected communities. The majority of recent deep learning methods follow a two-stage pipeline that separates representation learning from the final clustering step. In the first stage, a GNN encoder learns node embeddings in a self-supervised manner, often using contrastive learning frameworks like Deep Graph Infomax (DGI) [52], DMGI [39], and NACC[49]. In the second stage, a standard clustering algorithm like K-means is applied to cluster these embeddings. This decoupled pipeline, however, creates a significant objective mismatch. The GNN is trained to optimize a surrogate objective, such as a contrastive loss, which is not a direct measure of the clustering quality. An embedding optimized for general representation quality is not guaranteed to be optimal for partitioning the graph into dense communities.
@@
To resolve this mismatch, an end-to-end paradigm has emerged that directly optimizes a clustering-specific objective function. For single-layer networks, the recent work [50] introduced a fully differentiable version of the classic modularity metric, allowing a GNN to be trained to directly maximize the clustering quality. Our proposed LDGA framework extends this powerful insight to multi-layer networks. By developing a differentiable multi-layer modularity loss, we align the entire learning process with the goal of community detection in multilayer networks, unifying feature learning and graph clustering into an end-to-end trainable system.
@@ IV-A Node Feature Initialization
Our framework supports multilayer graphs with optional node attributes. If raw node features are provided, they can be directly incorporated as input to each layer. If no attributes are available, which is common in many benchmark datasets, we initialize features via an unsupervised embedding method such as DeepWalk [40], applied independently to each layer. This produces per-layer embeddings ${\bm{Z}}^{(s)}=\left[{\bm{z}}^{(s)}_{1},{\bm{z}}^{(s)}_{2},\dots,{\bm{z}}^{(s)}_{n}\right]^{T}\in\mathbb{R}^{N\times d}$ that capture local connectivity patterns while preserving layer identity. Stacking them yields a feature tensor ${\bm{Z}}\in\mathbb{R}^{L\times N\times d}$.
@@ IV-B Community-Latent Encoder (CLE)
Most existing deep learning methods focus on designing neural network architectures from the perspective of nodes and network layers [48, 54, 34], while overlooking community-level semantics. To enrich representations, we introduce a community-latent encoder (CLE), which enriches node features with learnable prototypes that represent latent communities. For each layer $s$, CLE maintains a learnable prototype matrix ${\bm{E}}^{(s)}\in\mathbb{R}^{\kappa\times d}$, where each row vector represents a learnable latent prototype for one community.
@@ IV-E Multi-layer Modularity Maximization Loss
Inspired by Tsitsulin et al.[50], who introduced a differentiable relaxation of the Newman–Girvan modularity, we extend this idea to multilayer network. We propose a loss function, ${\mathcal{L}}_{\text{soft}}$, which couples differentiable multilayer modularity with a normalized cluster balance regularizer:
@@ V-A Experiment Setup
To evaluate the performance of our proposed method, we conduct experiments on both synthetic and real-world datasets. Synthetic datasets mainly refer to mLFR [7], which is a baseline model for generating multilayer networks. The mixing parameter $\mu$ in the mLFR datasets, indicating the probability of edges connecting to the nodes of the other communities, can be adjusted. A smaller $\mu$ indicates more obvious community structures, while a larger $\mu$ indicates more vague community structures. Table I presents the detailed parameter settings of mLFR datasets in our experiment.
We also select ten extensive real-world multilayer networks datasets from domains covering society (AUCs [35], CKM [15], Lazega [31, 43], RM [20], Vickers [53]), biology (C.Elegans [11], Drosophila [46, 18], SACCHPOMB [46, 18]), finance Kapferer [27]), transportation (EUAir [10]). They are all collected manually and without ground-truth labels. Table II presents their basic information.
@@
In order to verify the performance of our method, we select nine state-of-the-art community detection methods, which can be divided into the following categories. The first group is traditional community detection algorithms: GenLouvain(GL) [37], M-Infomap(MI) [17], MDLPA [6], MolTi [19]. The second group is five multiplex embedding algorithms: HDMI [26], DMGI [39], GEAM [54], NFACC [47], NACC [49].
@@
We use common evaluation metrics of community detection to evaluate the performance of all the methods, including NMI (Normalized Mutual Information) [2], ARI (Adjusted Rand Index) [23], Purity [42], and multilayer modularity [38]. The maximum value of multilayer modularity is 1, and a higher modularity means a clearer community structure. NMI, ARI, and Purity are only applied to synthetic datasets since they need ground-truth labels, while the modularity of detected communities is applied to both synthetic and real-world datasets. The resolution parameter and coupling parameter of modularity are both set to 1. For all experiments, we run 10 trials with different random seeds, and report the best results in Tables III, IV, and Fig. 3.